# AlphaZero Neural Scaling and Zipf’s Law: a Tale of Board Games and Power Laws

### v1

@@ B.2 Zipf’s law in non-uniformly-random games
Appendix C Policy degradation with temperature
In section 5.2 we plot the correlation between the scaling-law and Zipf’s-law exponents by modulating the temperature of the action-selection policy, defined in Eq. 5. We claim that this correlation, plotted in Fig. 2B, is caused solely by the shifting Zipf distribution that affects agent performance, and not by temperature causing agents to play suboptimal moves.
@@ Appendix C Policy degradation with temperature
The set of optimal actions $A(s)$ is calculated with a Connect Four solver using alpha-beta pruning. An optimal action is defined as any action that, under optimal play, will lead to a victory, or to a draw if a victory is not possible. States where every move leads to a loss do not have any optimal actions, and are ignored in this analysis. The probability to ”blunder”, i.e. lose the lead in a game, is therefore equivalent to $1-\bm{p}(optimal)$.
@@
Figure 12: A zoomed-out version of Fig. 2C that includes higher temperatures ($=$ smaller exponents). Fig. 2C only contains data to the right of the vertical cutoff line, which marks the temperature where $T$ stops affecting policy quality (equivalent to the line in Fig. 11). This line happens to coincide with the point where the Zipf exponent starts changing.
Appendix D The difference between AlphaZero value and actor-critic value
In sections 6 and 7 we present results on AlphaZero loss scaling, focusing on the loss of the network’s value head. We chose the value head since it directly dictates the agent’s policy both at training and inference time. This appendix, aimed for readers unfamiliar with AlphaZero, clarifies how AlphaZero’s value head differs from the value net of actor-critic methods, which has no effect on policy at inference time.
Actor-critic methods are a family of popular policy-gradient algorithms, such as PPO and SAC (Konda & Tsitsiklis, 1999; Schulman et al., 2017; Haarnoja et al., 2018). These models use an architecture where two neural network outputs are used, namely a value estimation $v$ (or alternatively a $Q$ value) and a policy vector $\bm{p}$. Actions are sampled from the policy vector, which in turn is trained on the output of the value net $v$. While the loss of the value net on test data should be low for better agents, it has no effect on the actions played at test time.
@@ Appendix D The difference between AlphaZero value and actor-critic value
The visit count $N(s,a)$ is the number of times action $a$ was probed during MCTS, sampled in this equation for each action at the root node, meaning at the current game state $s_{0}$. The visit count is indirectly determined by the neural net’s value and policy outputs $(v,\bm{p})$, since these parameters are used by MCTS to decide which future actions to probe when performing rollouts. The decision what action to probe at each junction in the game tree is done by maximizing the following quantity (Silver et al., 2017b):
[…]
@@
The exploration term $U(s,a)$ is a variation of the PUCT algorithm (Rosin, 2011), favoring actions with low visit counts $N$:
@@
Appendix E Board position visualization
@@
Figure 13: Checkers board positions. Low-rank ($=$ high-frequency) states are mostly opening moves, visited shortly after the game begins. These states are mixed with high-frequency end-game positions, where only about 2-3 pieces still remain on the board, all promoted from man to king. The time it takes to end the game after reaching such an empty-board state is long; about half of all games end in a draw after no capture occurred for 40 turns.
Figure 14: Oware board positions. States below rank 100 are all openings, deviating only slightly from the initial position of 4 seeds in each pit. At rank $164$, end-game states start to dominate the distribution, making up $>80\%$ of all states. Most seeds have been captured in these late-game states, and the game ends shortly after they are played, when the active player has no seeds left on their row.
In Figs. 13 and 14 we plot randomly-sampled game states from the state-frequency distribution, for both Checkers and Oware games. High-frequency states split into two clusters, namely opening moves and end-game moves. Early-game openings, plotted in the top row of each figure, are all small deviations from the initial board state. In the initial state of Checkers, each players’ man pieces occupy their 3 closest rows; in the Oware initial state, there are 4 seeds in each pit.
@@
Appendix F Frequency scaling on a different axis: capture difference
@@
Figure 15: State frequency distribution by capture difference. Measuring the score difference between players provides another natural axis along which frequency drops smoothly. Agents could potentially prioritize optimizing small-difference states over large-difference, modelling more states with increasing capacity. Prioritizing states in this way is not equivalent to prioritizing states by their frequency-rank, i.e. by the Zipf’s law of Fig. 1, since a large portion of the states come from the tail-end of the Zipf distribution which is discarded in the main section results. If we ignore the tail of Fig. 1 by plotting only high-frequency states, the frequency distribution changes significantly (yellow). The actual order in which data is learned can probably not be visualized by a simple metric.

### v2

@@ C.2 Zipf’s law in non-uniformly-random games
Appendix D Policy degradation with temperature
In section 6 we plot the correlation between the scaling-law and Zipf’s-law exponents by modulating the temperature of the action-selection policy, defined in Eq. 5. We claim that this correlation, plotted in Fig. 2B, is caused solely by the shifting Zipf distribution that affects agent performance, and not by temperature causing agents to play suboptimal moves.
@@ Appendix D Policy degradation with temperature
The set of optimal actions $A(s)$ is calculated with a Connect Four solver using alpha-beta pruning. An optimal action is defined as any action that, under optimal play, will lead to a victory, or to a draw if a victory is not possible. States where every move leads to a loss do not have any optimal actions, and are ignored in this analysis. The probability to "blunder", i.e. lose the lead in a game, is therefore equivalent to $1-\bm{p}(optimal)$.
@@
Figure 12: A zoomed-out version of Fig. 2C that includes higher temperatures (== smaller exponents). Fig. 2C only contains data to the right of the vertical cutoff line, which marks the temperature where $T$ stops affecting policy quality (equivalent to the line in Fig. 11). This line happens to coincide with the point where the Zipf exponent starts changing.
Appendix E Ground-truth loss scaling on different distributions
We present here another plot of ground-truth value-loss scaling, in addition to Fig. 4B. In Fig. 4B, we plot the loss of agents on a collective dataset, aggregated from the training runs of several agents of varying sizes. Here in Fig. 13A we plot the same loss, this time by calculating each agent’s loss on their own dataset. The only difference between the two plots is the order of states; the ground-truth labels of each state remain the same. The error bars of Fig. 4 and Fig. 13 are the standard-deviation of agent loss across agents, either geometric when the y-axis is in logscale (Fig. 4A) or arithmetic when the y-axis is linear.
Figure 13: Connect Four ground-truth value-loss on different datasets. A: Average loss on each agent’s own training data. B: Fig. 4B, presented here again for comparison, displaying average loss on a common dataset collected by different agents.
Appendix F The difference between AlphaZero value and actor-critic value
In sections 7 and 8 we present results on AlphaZero loss scaling, focusing on the loss of the network’s value head. We chose the value head since it directly dictates the agent’s policy both at training and inference time. This appendix, aimed for readers unfamiliar with AlphaZero, clarifies how AlphaZero’s value head differs from the value net of actor-critic methods, which has no effect on policy at inference time.
Actor-critic methods are a family of popular policy-gradient algorithms, such as PPO and SAC [50, 51, 52]. These models use an architecture where two neural network outputs are used, namely a value estimation $v$ (or alternatively a $Q$ value) and a policy vector $\bm{p}$. Actions are sampled from the policy vector, which in turn is trained on the output of the value net $v$. While the loss of the value net on test data should be low for better agents, it has no effect on the actions played at test time.
@@ Appendix F The difference between AlphaZero value and actor-critic value
The visit count $N(s,a)$ is the number of times action $a$ was probed during MCTS, sampled in this equation for each action at the root node, meaning at the current game state $s_{0}$. The visit count is indirectly determined by the neural net’s value and policy outputs $(v,\bm{p})$, since these parameters are used by MCTS to decide which future actions to probe when performing rollouts. The decision what action to probe at each junction in the game tree is done by maximizing the following quantity [53]:
$\text{MCTS action}=\underset{a}{\mathrm{argmax}}\left(Q(s,a)+U(s,a)\right)$ (9)
@@
The exploration term $U(s,a)$ is a variation of the PUCT algorithm [54], favoring actions with low visit counts $N$:
@@
Appendix G Board position visualization
@@
Figure 14: Checkers board positions. Low-rank (== high-frequency) states are mostly opening moves, visited shortly after the game begins. These states are mixed with high-frequency end-game positions, where only about 2-3 pieces still remain on the board, all promoted from man to king. The time it takes to end the game after reaching such an empty-board state is long; about half of all games end in a draw after no capture occurred for 40 turns.
Figure 15: Oware board positions. States below rank 100 are all openings, deviating only slightly from the initial position of 4 seeds in each pit. At rank $164$, end-game states start to dominate the distribution, making up $80\%$ of all states. Most seeds have been captured in these late-game states, and the game ends shortly after they are played, when the active player has no seeds left on their row.
In Figs. 14 and 15 we plot randomly-sampled game states from the state-frequency distribution, for both Checkers and Oware games. High-frequency states split into two clusters, namely opening moves and end-game moves. Early-game openings, plotted in the top row of each figure, are all small deviations from the initial board state. In the initial state of Checkers, each players’ man pieces occupy their 3 closest rows; in the Oware initial state, there are 4 seeds in each pit.
@@
Appendix H Inverse scaling mechanism
In section 8.3, we mention a hypothetical mechanism that could cause the observed inverse scaling phenomenon. Here we discuss this mechanism in more detail.
We make the following main assumptions:
  • Agents fit states in descending order of frequency. An agent with a neural network of size $N$ can perfectly fit the values of $n=N/c$ states, where $c$ is the neural capacity needed to fit a single state [12].
  • As the agent improves during training, game outcomes (i.e. state values) change. The values of early-game states change significantly during training, while the values of late-game states change only slightly, if at all.
  • The shift of state values causes neural plasticity loss due to dormant neurons [44].
The first assumption is based on the quantization model of neural scaling [12].
The second assumption is based on the fact that, in most board games, player skill greatly influences the value of early-game states. In the games discussed in this paper, a match between random agents will usually end in a draw ($v=0$) while a match between optimal agents will end in a first-player victory ($v=\pm 1$), except for checkers. In contrast, late game states that are played a few steps before the game ends have values that are largely independent of player skill, due to power imbalance. Very late states have fixed values if MCTS can probe the entire game tree stemming from them.
The third assumption is based on the work of Sokar et al. (2023) [44], who showed that changing the labels of a dataset during training causes some neuron activations to die out. These so-called "dormant neurons" effectively reduce the neural capacity of the model.
Assume we train a small agent with $10^{m}$ neurons and a large agent with $10^{k}$ neurons, $k>m$. In our model, the small agent roughly fits $10^{m}/c\approx n_{early}$ states, where $n_{early}$ are the number of early-game states with higher frequency than late game states. In Oware and Checkers, $n_{early}$ falls in the range $[100,200]$, see Fig. 6. The large agent fits $10^{k}/c\gg n_{early}$ and mostly memorizes the values of late-game states.
For simplicity, let us treat the change of state values as a single, abrupt event during training. After this event, both models have high loss on early-game states, but maintain the same loss as before on late-game states, since the latter group kept their value labels. Due to plasticity loss, both models can now fit fewer states than before. The small agent now has high loss, since it could only fit the first $n_{early}$ states before values changed, meaning most of the states it memorized now have new labels. The large model, on the other hand, maintains a lower loss, since it could fit orders of magnitude more states than the first $n_{early}$ states, most of which are late-game states.
It is likely that the small agent, which now has loss close to that of a random agent, will continue to train and re-learn as many states as it can, starting from the most frequent ones. Eventually it will regain low loss on most or all of the first $n_{early}$ states. In contrast, the large agent maintains low loss, and could be close to a local minimum of the loss landscape that fits late-game states well but assigns wrong labels to early-game states. To leave this minimum and re-learn the first $n_{early}$ states, the large agent will have to sacrifice knowledge of some late-game states, since it lost a fraction of its neural capacity. In this scenario, the large model could remain stuck in the local loss minimum and never regain performance on early-game states.
In reality, it is more likely that value labels shift gradually rather than abruptly, meaning that loss increases gradually during training rather than spiking abruptly. We observe such a loss degradation in Fig. 8, where large agent value loss on early-game states starts to degrade mid-training.
Appendix I Frequency scaling on a different axis: capture difference
@@
Figure 16: State frequency distribution by capture difference. Measuring the score difference between players provides another natural axis along which frequency drops smoothly. Agents could potentially prioritize optimizing small-difference states over large-difference, modelling more states with increasing capacity. Prioritizing states in this way is not equivalent to prioritizing states by their frequency-rank, i.e. by the Zipf’s law of Fig. 1, since a large portion of the states come from the tail-end of the Zipf distribution which is discarded in the main section results. If we ignore the tail of Fig. 1 by plotting only high-frequency states, the frequency distribution changes significantly (yellow). It is likely impossible to visualize the true order in which data is learned with a simple metric.
Appendix J Distribution of ideal games
In section 5 we claim that random games played in a toy-model setting will result in a frequency distribution shaped as a series of plateaus, centered around a power law of exponent $\alpha=1$. We verify this claim here.
In short, in a simple branching game the number of possible states is exponential in the turn number $t$, and all states at turn $t$ have an equal probability to appear in a game, hence the series of frequency plateaus observed in Fig. 2A.
Frequency distribution function
Consider a random game setting following the rules presented in section 5, namely:
  • Each game lasts $K$ turns.
  • Each turn $t$, a move $a_{t}$ is sampled uniformly from $b$ options.
  • Each board position $s$ played at turn $t(s)$ can only be reached by a single, unique sequence of moves $\bm{a}_{s}=\left(a_{1},a_{2},\dots a_{t(s)}\right)$. This means that one can define a state $s$ only by the sequence of moves played to reach it.
The position at turn $t$ is effectively sampled uniformly from $b^{t}$ possible positions, since the $t$ moves $a_{i}$, $i\in\{1,\dots,t\}$ that lead to it are each sampled uniformly from $b$ options. Since $K$ turns are played each game, when sampling a board position $S$ randomly from all games played the frequency of sampling board state $s$ is:
$P(S=s)=P(t=t(s))\cdot P(S=s|t=t(s))=\frac{1}{K}\cdot\frac{1}{b^{t}}\,$, (13)
where $t(s)$ is the turn number of board state $s$ and $P(t=t(s))$ is the probability to sample from turn $t(s)$.
Sorting the board positions in descending order of frequency results in a series of plateaus, each corresponding to a turn number. The number of board states in plateau number $t$ is $b^{t}$. Plateau number $t$ starts after board position number $n_{start}(t)$, defined by:
$n_{start}(t)=\sum_{i=1}^{t-1}b^{i}=\frac{b^{t}-b}{b-1}\,$, (14)
since each preceding plateau $i$ contains $b^{i}$ states. Board state number $n$ belongs to plateau number $t$ for the integer $t\in\mathbb{N}$ that satisfies:
$n_{start}(t)\leq n<n_{start}(t+1)\,$, (15)
or:
$\frac{b^{t}-b}{b-1}\leq n<\frac{b^{t+1}-b}{b-1}\,.$ (16)
Shifting terms in the inequality and taking the log we get:
$b^{t}-b$ $\leq{}$ $(b-1)n$ $<b^{t+1}-b$ (17)
$b^{t}$ $\leq{}$ $(b-1)n+b$ $<b^{t+1}$ (18)
$t\log(b)$ $\leq{}$ $\log\left[(b-1)n+b\right]$ $<(t+1)\log(b)$ (19)
$t$ $\leq{}$ $\frac{\log\left[(b-1)n+b\right]}{\log b}$ $<t+1$ (20)
since $t\in\mathbb{N}$, this is equivalent to using the floor operator:
$t(n)=\left\lfloor\frac{\log\left[(b-1)n+b\right]}{\log b}\right\rfloor\,.$ (21)
Using Eq. 13 one sees that the frequency of the $n$-th most common board state, $s_{n}$, is proportional to:
$P(S=s_{n})=\frac{1}{K}\cdot\frac{1}{b^{t(n)}}\,$, (22)
where $t(n)$ is defined in Eq. 21. This is the series of plateaus seen in Fig. 2A.
Distribution is bounded around power law
Let us define $\tilde{t}(n)\in\mathbb{R}$ as $t(n)$ without the floor operator:
$\tilde{t}(n)=\frac{\log\left[(b-1)n+b\right]}{\log b}\,.$ (23)
Since $x-1<\left\lfloor x\right\rfloor\leq x$, $P(S=s_{n})$ is bounded by:
$\frac{1}{K}\frac{1}{b^{\tilde{t}(n)}}\leq P(S=s_{n})<\frac{1}{K}\frac{1}{b^{\tilde{t}(n)-1}}\,.$ (24)
Expanding the bounding functions we get:
$\frac{1}{b^{\tilde{t}(n)}}=\frac{1}{b^{\frac{\log\left[(b-1)n+b\right]}{\log b}}}=\frac{1}{e^{\log\left[(b-1)n+b\right]}}=\frac{1}{(b-1)n+b}\,$, (25)
leaving us with:
$\frac{1}{K}\frac{1}{(b-1)n+b}\leq P(S=s_{n})<\frac{1}{K}\frac{b}{(b-1)n+b}\,.$ (26)
We therefore see that the frequency distribution $P(S=s_{n})$ is bounded by two straight lines (power laws with power 1) with different coefficients, which appear as two parallel lines in log-log scale, as we observe in Fig. 2A.